Question 1.2 [2, 2, 2]


Suppose that we have two events A and B such that P A( ) 0.8 = and P B( ) 0.7 = .


a) Is it possible that PA B ( ) 0.1 ∩ = ? Explain your answer.


b) What is the smallest possible value of PA B ( ) ∩ ?


c) What is the largest possible value of PA B ( ) ∩ ?

1
Expert's answer
2022-06-07T12:20:36-0400
"P(A\cup B)=P(A)+P(B)-P(A\cap B)"

"P(A\cap B)=P(A)+P(B)-P(A\cup B)"

Since "P(A)=0.8, P(B)=0.7,0\le P(A\cup B)\le1,"

"P(A\cap B)\le \min{(P(A), P(B))}" then


"P(A\cup B)=P(A)+P(B)-P(A\cap B)"


"0\le P(A)+P(B)-P(A\cap B)\le 1"

"0\le 0.8+0.7-P(A\cup B)\le 1"

"P(A\cap B)\le P(B), P(A\cap B)\ge 08+0.7-1"


"0.5\le P(A\cap B)\le 0.7"

a) It is impossible that "P(A\cap B)=0.1."


b) The smallest possible value of "P(A\cap B)" is "0.5."


c) The largest possible value of "P(A\cap B)" is "0.7."